# ECM: A Unified Electronic Circuit Model for Explaining the Emergence of In-Context Learning and Chain-of-Thought in Large Language Model

### v1

@@ 5.1 ICL Exploration
Overview of the fine-grained self-consistency strategy in circuits. By decomposing parallel CoT resistors into smaller units, this approach effectively reduces overall circuit resistance value. b, Overview of the Chain-of-Verfication (CoV) strategy in circuits. In addition to optimizing parallel CoT resistances, this method incorporates the concept of voting resistors in self-consistency to further reduce overall circuit resistance. The CoV strategy achieves this by sampling multiple voting results and validating them to finalize the decision. c, Leveraging stronger magnetic fields and reducing logical resistance significantly enhances circuit performance. The power-enhanced model achieved notable results in the IOI (left figure) and the IMO (right figure). d, The methodology for incorporating sufficient theoretical power into manual ratings during idea generation scenarios. This approach demonstrates a significant improvement over the original solution by enhancing the accuracy and reliability of the evaluation process.
5.2 CoT Exploration
Optimize the same proportion of capabilities. The greater the original difficulty of the task corresponding to the capability, the greater the impact on the final performance.
We can prove that optimizing a resistor with higher resistance to its original value $\frac{1}{k}$ results in a greater reduction in overall resistance of reasoning resistor $R_{\text{CoT}}$, yielding a more significant performance improvement compared to optimizing a lower resistance (See in Theoretical Proof). As shown in the meta-analysis in Appendix A, when solving questions that demand advanced planning, o1 [28]’s reasoning method, which emphasizes extensive exploration, significantly enhances the most challenging reasoning components. This improvement ultimately boosts overall performance on complex reasoning tasks. However, for tasks where the hardest component requires extensive domain-specific knowledge, insufficient optimization of such knowledge limits performance gains in those domains.
Fine-grained Self-consistency achieves better performance
@@ 5.2 CoT Exploration
Easily, we can prove that, with a stable supply voltage, we can split a self-consistency ensemble into more fine-grained step-by-step self-consistency by conducting majority voting verification for each step, so that each resistance value of verification resistor can be smaller, and we can obtain larger output power (see Theoretical Proof). Indeed, as the number of self-consistency steps increases, the theoretical power value gradually increases. As shown in Figure 6a, the corresponding performance also gradually increases. At the same time, the corresponding upper bound and lower bound also satisfy the original self-consistency and coverage. This not only proves the effectiveness of our proposed fine-grained self-consistency method but also proves the practical usability of the theoretical model.
Chain-of-Verification enhances model performance with parallelized verification resistors
In conventional self-consistency frameworks, verification is typically conducted through series resistors. We propose an innovative approach, referred to as the “chain-of-verification,” which involves two key looped stages: first, verifying the correctness of individual outputs in parallel; and second, merging these verified results and subsequently meta-validating the correct to confirm the overall result of the merged outputs. It can enhance model performance by reducing the overall verification resistance. As depicted in Figure 6b, parallelization initially boosts performance by distributing verification tasks and lowering verification resistance. However, this improvement has limits. As parallelization increases, the overall resistors for parallel verifiers begins to rise. Beyond a critical threshold, this combined resistor surpasses the original self-consistency resistor, reversing the performance gains. As depicted in Figure 6b, this also demonstrates excessive chain-of-verification introduces diminishing returns. Thus, the chain-of-verification framework underscores the importance of balancing parallelization and resistor to optimize model verification efficiency.
6 Best Practice
To evaluate the impact of sufficient power on LLM performance, we analyzed its effects across three domains: the International Olympiad in Informatics (IOI), the International Mathematical Olympiad (IMO), and Natural Language Processing (NLP) idea generation using the AI-Scientist framework [23], by ensembling strategies outlined in Exploration Section. Our findings, shown in Figure 6c, demonstrate that augmenting the o1-preview model with enhanced power results in significant improvements. The power-enhanced o1-preview outperformed AlphaCode [20] and other state-of-the-art models on the IOI-level test set (code_contests [20]). Crucially, this boost consistently surpassed configurations lacking sufficient power, underscoring its pivotal role in model efficacy. In competitive contexts, such as IOI 2024 and IMO 2024, the powered o1-preview outperformed nearly 80% of global participants, earning silver and bronze medals. This highlights power as a decisive factor in high-stakes scenarios where traditional optimizations often fall short. Beyond correctness-focused tasks, as shown in Figure 6d, sufficient power also enhanced performance in creative challenges. For instance, in NLP-related idea-generation, it facilitates the better generation of novel and impactful research ideas (over 1 point of overall score improvement on average), expanding the potential of LLMs in both objective and subjective tasks.
@@ 7 Conclusion & Discussion
The Electronic Circuit Model (ECM) offers a unified framework for understanding and optimizing the ICL and CoT mechanisms, marking a significant advancement in LLM research. This framework elucidates model behavior while providing a quantifiable approach to enhance LLM performance. Our experiments demonstrate that ECM effectively captures model intricacies, serving as a reliable tool for analyzing and improving AI capabilities. Moreover, applying power-based optimization derived from ECM to advanced reasoning challenges, such as the International Olympiad in Informatics and the International Mathematical Olympiad, yielded superior results compared to leading competitors. These outcomes validate the theoretical foundation of ECM and highlight its transformative potential in improving AI reasoning and learning efficiency, with broad implications for next-generation intelligent systems.
@@ 8 Code Availability
The code and relevant data are available at https://github.com/LightChen233/ECM.
A Method Implementation Details
A.1 Theoretical Model
Faraday’s Law of In-Context Learning
In-context learning (ICL) can be likened to the behavior of a semantic magnetic field generated by a sub-power source, as described by Faraday’s Law. The positive direction of the semantic magnetic field is defined by the query semantic vector $\overrightarrow{S}_{q}$. Given $N$ context samples represented by semantic vectors $\overrightarrow{S}_{i}$, the total semantic magnetic field strength $\Phi^{B}_{0}$ experienced by the model is the projection of these vectors onto $\overrightarrow{S}_{q}$. Mathematically:
@@ A.1 Theoretical Model
where $\lambda$ is the decay constant representing the rate at which the model forgets or absorbs information. According to Faraday’s Law of electromagnetic induction, the rate of change in the semantic magnetic field induces an electromotive force, which drives effective ICL. With a time step $\Delta t$ for each reasoning step, the induced EMF, $\mathcal{E}_{\text{ICL}}$, is:
$\mathcal{E}_{\text{ICL}}=-\frac{\Delta\Phi^{B}(t)}{\Delta t}=\lambda\Phi^{B}_{0}.$ (6)
@@
$\mathcal{E}_{\text{ICL}}=\lambda\sum^{N}_{i=1}\frac{\overrightarrow{S}_{q}\cdot\overrightarrow{S}_{i}}{\left\lvert\overrightarrow{S}_{q}\right\rvert}.$ (7)
In conclusion, this analogy offers a structured framework to analyze how information flows during ICL. By extending the dynamics of ICL through the principles of electromagnetic induction, we provide a mathematical model to explain how models incorporate and gradually forget information as they process contextual examples.
Ohm’s Law of Chain-of-Thought
In this framework, each sub-difficulty of a model introduces a distinct level of difficulty, analogous to resistors in series. Formally, following Chen et al. [5], the ease with which a model can solve a particular task is represented by its reasoning boundary. Let $\mathcal{B}_{\text{CoT}}$ denote the overall reasoning boundary of a Chain-of-Thought process, and $\mathcal{B}_{i}$ denote the reasoning boundary of the $i$-th sub-capability. The overall reasoning boundary can be expressed by the harmonic sum:
$\mathcal{B}_{\text{CoT}}=\frac{1}{\sum^{K}_{i=1}\frac{1}{\mathcal{B}_{i}}},$ (8)
where $K$ denotes the number of sub-capabilities. Reasoning resistor, $R$, quantifies the difficulty of a reasoning process and is mathematically defined as the reciprocal of the reasoning boundary. A higher boundary corresponds to an easier problem and lower resistor. Specifically, we define the sub-difficulty of the $i$-th sub-task as:
@@
Substituting this into Equation 8, the total reasoning resistor for the CoT process is expressed as:
$R_{\text{CoT}}=\sum^{K}_{i=1}R_{i},$ (10)
@@
where $R_{\text{CoT}}$ represents the combined difficulty, analogous to total resistor in a series circuit. Overall, the CoT reasoning framework is further formalized by the Ohm’s law analogy:
$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ICL}}}{R_{\text{CoT}}+R_{0}},$ (11)
@@
where $R_{0}$ denotes a output resistor with static resistance value. This modular formulation effectively quantifies cognitive task complexity through combined reasoning resistor.
A.2 Genral Experimental Settings
@@
All experiments, unless stated otherwise, are conducted on GPT-3.5-Turbo. Following Wei et al. [41], CoT experiments select 5 manually constructed demonstrations for all multi-step reasoning tasks. Top-p is selected from $\{0.95,1\}$, and Temperature is selected in range $[0,1]$ for robustness evaluation. Experiments are primarily conducted on the BigGSM dataset [5], a benchmark for mathematical reasoning. In ICL, demonstrations are dynamically retrieved using a 5-shot approach from the GSM8K development dataset [8] by default. To formalize the demonstration selection process, we retrieve the semantic magnetic field strength $\Phi^{B}_{0}$ by default, as defined in Eq. 4. All demonstration representation methods employ the meta-recognition [10] framework by default: GPT-4o [1] extracts sub-capability tags, and Roberta [22] concatenates and represents these tags in the hidden space. In addition, only the magnetic field strength $\Phi^{B}_{0}$ is calculated and the resistor value $R_{\text{CoT}}$ directly adopts the fitting values from Chen et al. [5]. In addition, other parameters are fitted values using the verification set. Different models fit different voltages $\mathcal{E}_{model}$. The whole experiment shares a $R_{0}$ value. Different demonstration representation methods have different $\lambda$. All experiments involving variance were sampled 3 times at different temperatures.
A.3 Model Verification Experiment Details
CoT Ohm’s Law Verification
@@ A.3 Model Verification Experiment Details
To validate the Ohm’s Law assumption in the CoT approach, we conduct the following experiment. To eliminate potential interference from ICL and ensure effective stimulation of CoT capabilities, we append the phrase “Let’s think step-by-step!" to the user query. This adjustment aims to stimulate and isolate the CoT mechanism and prevent ICL-related influences. We then perform sampling based on power levels from the BigGSM dataset, using 30-unit intervals to map calculated power to accuracy [^1] The absolute power values are meaningless due to inconsistencies in semantic space scaling across encoding methods. However, their relative values within a method correlate positively with real accuracy.. To reduce random errors and ensure sufficient sampling points, we retain data points with at least 10 samples. Accuracy is calculated for each sampling interval to analyze the relationship between power levels and model performance, enabling an effective evaluation of the correlation between sampling power and accuracy.
ICL Faraday’s Law Verification
@@
To validate the semantic magnetic field hypothesis in ICL, we observe that without CoT reasoning, performance on complex mathematical tasks often declines to near-zero levels, thereby cannot easily observe their relevance. Thus, this experiment incorporates both CoT reasoning and ICL. While the CoT reasoning format remains constant, we focus on varying the similarity measures used for sample retrieval through ICL. The central adjustment involves applying different similarity measures for sample retrieval, directly affecting the model’s performance on complex tasks. These measures also help calculate the additional “voltage" introduced by semantic magnetic fields, quantifying their influence on performance. All demonstration representation methods employ BGE [4] to represent the request in the hidden space.
Unified circuit mechanism verification
To further validate the effectiveness of our unified framework for modeling ICL and CoT, we conduct experiments integrating both paradigms into different retrieval settings. These experiments test the robustness and rationality of the proposed theoretical framework. Specifically, we examine three retrieval methods: random retrieval, forward magnetic field retrieval, and reverse magnetic field retrieval. In the random retrieval setup, samples are selected randomly from the development set as ICL demonstrations for each test sample. The Retrieval-ICL selects samples with maximized magnetic field intensity in the corresponding test sample’s direction. Conversely, the Reverse-ICL focuses on samples with minimal (even negative) magnetic field intensity, enabling exploration of contrasting ICL examples. This comprehensive validation highlights the robustness of our framework across diverse retrieval strategies and empirically demonstrates its generalizability in both ICL and CoT scenarios.
Extension Experiment Details on Various Models and Tasks
To further enhance the applicability of ECM, we conducted experiments across a series of models, where each model was designed to fit specific baseline voltage values, $\Phi^{B}_{0}$, and compute the corresponding power output. Subsequently, all models were consolidated onto a single graph, enabling us to analyze the correlation between output power and relative accuracy. As illustrated in Figure x, the results reveal clear and detailed correlations across the models. However, when accuracy exceeds 90%, the scarcity of samples requiring sufficient power in this range, likely due to limitations in the dataset, leads to poor fitting in this region.
Furthermore, our experiments on HotpotQA and Medical Probing closely adhered to the experimental settings established by Chen et al. [5]. However, since Medical Probing includes rationale annotations limited to background knowledge rather than chain reasoning, we supplemented portions of the reasoning process to facilitate a more precise correlation analysis. Additionally, the data for robotic planning was directly extracted from the running logs provided by Valmeekam et al. [34].
A.4 Model Explanation Experiments
A.4.1 ICL Explanation
Embedder Comparison

### v2

@@ 3.3 Exploring Novel Optimization Strategies through Magnetic Field Adjustment
a, Overview of the fine-grained self-consistency strategy in circuits. By decomposing parallel ITR resistors into smaller units, this approach effectively reduces overall circuit resistance. b, The Chain-of-Verification (CoV) strategy further optimizes circuit resistance by incorporating voting resistors into the self-consistency framework. This method reduces overall resistance by sampling and validating multiple voting results. c, Enhanced circuit performance is achieved by leveraging stronger magnetic fields and reducing logical resistance. The power-optimized model produced notable results in the IOI (left) and the IMO (right) tests. d, A methodology for integrating theoretical power into manual ratings during idea generation. This approach significantly improves the accuracy and reliability of the evaluation process compared to the original solution.
3.4 Exploring Novel Optimization Strategies through Resistor Reorganization
Optimising an equal proportion of model capabilities yields a performance gain that scales with the intrinsic difficulty of each capability.
Reducing the resistance of the most challenging reasoning component from $R$ to $\frac{R}{k}$ therefore produces a larger decrease in the total inference-time resistance, $R_{\text{ITR}}$ , than an equivalent change applied to an easier component (See Theoretical Proof in Appendix for mathematical proof). Meta-analysis (Methods) of benchmark problems requiring long-horizon planning confirms this prediction: the exploration-centred reasoning procedure of model o1-preview [31] preferentially improves high-resistance components and delivers the largest gains on the most complex tasks. By contrast, tasks dominated by specialised domain knowledge benefit less, owing to limited optimisation of the corresponding knowledge component.
Fine-grained self-consistency reduces verification resistance.
@@ 3.4 Exploring Novel Optimization Strategies through Resistor Reorganization
Easily, we can prove that, with a stable supply voltage, we can split a self-consistency ensemble into more fine-grained step-by-step self-consistency by conducting majority voting verification for each step, so that each resistance value of verification resistor can be smaller, and we can obtain larger output power (see Theoretical Proof). Indeed, as the number of self-consistency steps increases, the theoretical power value gradually increases. As shown in Fig. 7a, the corresponding performance also gradually increases. At the same time, the corresponding upper bound and lower bound also satisfy the original self-consistency and coverage. This not only proves the effectiveness of our proposed fine-grained self-consistency method but also proves the practical usability of the theoretical model.
Chain-of-verification balances parallel and serial checks.
Traditional self-consistency frameworks rely on a sequential verification, where each step depends on the completion of the previous one. This structure often incurs high computational costs, as subsequent tasks cannot begin until earlier ones are resolved. To overcome this limitation, we introduce the “chain-of-verification (CoV)," a novel strategy that parallelizes verification tasks into independent components. Each component operates concurrently, validating distinct parts of the system without interdependence. This parallel approach reduces overall computational resistance by distributing the workload across multiple verifiers, thereby enabling more effective validation. As illustrated in Fig. 7b, parallelization not only accelerates the verification process but also enhances model performance by mitigating bottlenecks. This improvement is particularly pronounced in large-scale systems, where sequential verification would otherwise impede efficiency. The effectiveness of CoV can also be mathematically proved in Theoretical Proof.
3.5 Unified Optimization for Downstream Applications
Adequate computational power is a primary determinant of LLM performance across disparate tasks. We benchmarked an enhanced version of the o1-preview model on algorithmic problem solving (International Olympiad in Informatics, IOI), mathematical reasoning (International Mathematical Olympiad, IMO) and open-ended idea generation in the AI-Scientist framework [28], using the ensemble procedure described above.
Increasing the output power yielded marked gains (Figure 7c): on the IOI code_contests [24] test set the power-augmented model outperformed AlphaCode [24] and all other published baselines. In live competitions (IOI 2024 and IMO 2024) it surpassed 8̃0% of human participants, earning silver and bronze medals, respectively, and highlighting compute as a critical factor when conventional optimisation is insufficient.
Enhanced compute also improved performance in creativity-oriented tasks. In the AI-Scientist benchmark, the same configuration increased the mean expert-rated novelty and impact scores by >1.0 (Fig. 7d), indicating that scaling compute extends the applicability of LLMs beyond correctness-centred objectives.
@@ 4 Conclusion & Discussion
The Electronic Circuit Principles (ECP) introduces a unified principle for understanding and optimizing LLMs’ performance. By quantifying internal dynamics, ECP clarifies model behavior and guides modular performance improvements. Our experiments confirm that ECP reliably captures model complexities, providing a robust tool for AI analysis and enhancement. Moreover, ECP-driven power optimizations applied to the International Olympiad in Informatics and the International Mathematical Olympiad outperform current leading methods. These findings validate ECP’s theoretical foundation and underscore its potential to elevate AI reasoning and learning efficiency in next-generation systems.
@@ 5 Code Availability
The code and relevant data are available at https://github.com/LightChen233/ECP.
A Methods
A.1 Electronic Circuit Principles (ECP)
Faraday’s Law of Inference-Time Learning
Inference-time learning (ITL), also named as in-context learning, can be likened to the behavior of a semantic magnetic field generated by a sub-power source, as described by Faraday’s Law. The positive direction of the semantic magnetic field is defined by the query semantic vector $\overrightarrow{S}_{q}$. Given $N$ context samples represented by semantic vectors $\overrightarrow{S}_{i}$, the total semantic magnetic field strength $\Phi^{B}_{0}$ experienced by the model is the projection of these vectors onto $\overrightarrow{S}_{q}$. Mathematically:
@@ A.1 Electronic Circuit Principles (ECP)
where $\lambda$ is the decay constant representing the rate at which the model forgets or absorbs information. According to Faraday’s Law of electromagnetic induction, the rate of change in the semantic magnetic field induces an electromotive force, which drives effective ITL. With a time step $\Delta t$ for each reasoning step, the induced EMF, $\mathcal{E}_{\text{ITL}}$, is:
$\mathcal{E}_{\text{ITL}}=-\frac{\Delta\Phi^{B}(t)}{\Delta t}=\lambda\Phi^{B}_{0}.$ (7)
@@
$\mathcal{E}_{\text{ITL}}=\lambda\sum^{N}_{i=1}\frac{\overrightarrow{S}_{q}\cdot\overrightarrow{S}_{i}}{\left\lvert\overrightarrow{S}_{q}\right\rvert}$. (8)
In conclusion, this analogy offers a structured framework to analyze how information flows during ITL. By extending the dynamics of ITL through the principles of electromagnetic induction, we provide a mathematical principle to explain how models incorporate and gradually forget information as they process contextual examples.
Ohm’s Law of Inference-Time Reasoning
In this framework, each sub-difficulty of a model introduces a distinct level of difficulty, analogous to resistors in series. Formally, following Chen et al. [4], the ease with which a model can solve a particular task is represented by its reasoning boundary. Let $\mathcal{B}_{\text{ITR}}$ denote the overall reasoning boundary of a Inference-Time Reasoning process, and $\mathcal{B}_{i}$ denote the reasoning boundary of the $i$-th sub-capability. The overall reasoning boundary can be expressed by the harmonic sum:
$\mathcal{B}_{\text{ITR}}=\frac{1}{\sum^{K}_{i=1}\frac{1}{\mathcal{B}_{i}}},$ (9)
where $K$ denotes the number of sub-capabilities. Reasoning resistor, $R$, quantifies the difficulty of a reasoning process and is mathematically defined as the reciprocal of the reasoning boundary. A higher boundary corresponds to an easier problem and lower resistance. Specifically, we define the sub-difficulty of the $i$-th sub-task as:
@@
Substituting this into Equation 9, the total reasoning resistor for the ITR process is expressed as:
$R_{\text{ITR}}=\sum^{K}_{i=1}R_{i},$ (11)
@@
where $R_{\text{ITR}}$ represents the combined difficulty, analogous to a total resistor in a series circuit. Overall, the ITR reasoning framework is further formalized by the Ohm’s law analogy:
$I_{model}=\frac{\mathcal{E}_{\text{model}}+\mathcal{E}_{\text{ITL}}}{R_{\text{ITR}}+R_{0}},$ (12)
@@
where $R_{0}$ denotes an output resistor with static resistance value. This modular formulation effectively quantifies cognitive task complexity through a combined reasoning resistor.
A.2 General Experimental Settings
@@
All experiments, unless stated otherwise, are conducted on GPT-3.5-Turbo. Following Wei et al. [43], ITR experiments select 5 manually constructed demonstrations for all multi-step reasoning tasks. Top-p is selected from $\{0.95,1\}$, and Temperature is selected in range $[0,1]$ for robustness evaluation. To formalize the demonstration selection process, we retrieve the semantic magnetic field strength $\Phi^{B}_{0}$ by default, as defined in Eq. 5. In addition, only the magnetic field strength $\Phi^{B}_{0}$ is calculated and the resistor value $R_{\text{ITR}}$ directly adopts the fitting values from Chen et al. [4]. Further, other parameters are fitted values using the verification set. Different models fit different voltages $\mathcal{E}_{model}$. The whole experiment shares a $R_{0}$ value. Different demonstration representation methods have different $\lambda$. All experiments involving variance were sampled 3 times at different temperatures. Notably, in addition to the validation experiments, we predominantly utilize the BigGSM [4] and BigGSM++ [5] datasets, which precisely annotate reasoning difficulty across global planning, local operations, and calculation sub-tasks. In ITL, demonstrations are dynamically retrieved using a 5-shot approach from the GSM8K development dataset [8] by default.
A.3 Verification Experiment Details
In the effectiveness validation section of ECP, and with the aim of comprehensively assessing the efficacy of our tasks, we first classify each task’s difficulty (hereafter “resistance”) into four categories: global planning difficulty, local operational difficulty, domain‐knowledge analysis difficulty, and computational difficulty (the latter being relevant only to mathematical tasks). The total resistance is then defined as:
$R_{\text{ITR}}=R_{\text{plan}}+R_{\text{operation}}+R_{\text{domain}}+R_{\text{calculate}},$ (13)
where $R_{\text{plan}}$ denotes the resistance due to global planning difficulty, $R_{\text{operation}}$ denotes the resistance due to local operational difficulty, $R_{\text{domain}}$ denotes the resistance due to domain‐knowledge analysis difficulty, and $R_{\text{calculate}}$ denotes the resistance due to computational difficulty (set to zero for non‐computational tasks).
Furthermore, as illustrated in the Fig. 8, the task evaluation benchmarks are primarily drawn from the BigGSM [4], BigGSM++ [5], MedProbing [7], Med-MCQA [32], MedQA [18], PubMedQA [19], BBH [38], MMLU [16], MMLU-Pro [42], GPQA [36], and SuperGPQA [12] datasets with over 350 tasks.
[…]
The annotation of the number of global and local reasoning steps required depends on the model-generated correct rationales and is automated through a manually constructed step-segmentation annotation program. Furthermore, since the granularity of the model’s reasoning does not fully align with that of humans, and the segmentation program cannot perform fine-grained annotations for every case, a considerable amount of annotation noise exists in the validation experiments across nearly 350 tasks, resulting in a relatively low $R^{2}$. However, under a strictly human-annotated setting for global and local step counts, such as that of the BigGSM++ dataset, the $R^{2}$ can exceed $0.9$. Therefore, except for the validation experiments, all subsequent experiments are conducted on the BigGSM++ dataset.
Figure 8: Over 350 tasks in unified evaluation data.
Inspired by Chen et al. [4], we treat the domain difficulty as a constant for each task; this constant was obtained via parameter search using Newton’s method in a validation experiment on 10% of the GPT-3.5-turbo generated dataset. For mathematical tasks, we design task-specific regular-expression extraction rules to estimate computational difficulty; however, since most operations not involving large numbers yield a computational difficulty close to zero, we omit this factor for some math tasks.
ITR Ohm’s Law Verification
@@ A.3 Verification Experiment Details
To validate the Ohm’s Law assumption in the ITR approach, we conduct the following experiment. To eliminate potential interference from ITL and ensure effective stimulation of ITR capabilities, we append the phrase “Let’s think step-by-step!" to the user query. This adjustment aims to stimulate and isolate the ITR mechanism and prevent ITL-related influences. We then perform sampling based on power levels from the BigGSM dataset, using task-specific unit intervals to map calculated power to accuracy [^1] Since the distribution of reasoning difficulty varies across tasks, a sufficiently large sample size is required to accurately compute accuracy. Accordingly, we employ different units for each task.. To reduce random errors and ensure sufficient sampling points, we retain data points with at least 10 samples. Accuracy is calculated for each sampling interval to analyze the relationship between power levels and model performance, enabling an effective evaluation of the correlation between output power and accuracy to verify Ohm’s Law.
Figure 9: The Analysis for ITL Faraday’s Law Verification Across 350 Tasks and 4 Models.
ITL Faraday’s Law Verification
@@
To validate the semantic magnetic field hypothesis in ITL, we observe that without ITR reasoning, performance on complex mathematical tasks often declines to near-zero levels, thereby cannot easily observe their relevance. Thus, this experiment incorporates both ITR reasoning and ITL. While the ITR reasoning format remains constant, we focus on varying the similarity measures used for sample retrieval through ITL. The central adjustment involves applying different similarity measures for sample retrieval, directly affecting the model’s performance on complex tasks. These measures also help calculate the additional “voltage" introduced by semantic magnetic fields, quantifying their influence on performance. All demonstration representation methods employ BGE [3] to represent the request in the hidden space.
A.4 Optimization Interpreting Experiments through Magnetic Field Adjustment
Encoder Comparison